OpenStudy (anonymous):

Find the equation in standard form which passes through (–2, –7) and (1, –1). You must show all work.

OpenStudy (hoblos):

y=2x-3

OpenStudy (anonymous):

y=mx+c m=slope \[m=-1-(-7)/1-(-2)\] m=2 y=2x+c put any point in the equation lets put (1,-1) -1=2(1)+c c=-3 the equation is y=2x-3